Mrs. Levy sells cookie bouquets. She can make no more than 15 bouquets in one week. Small bouquets sell for $25 and large bouquets sell for $55. If she prefers to make no more than 7 large arrangements a week, how many of each size should she sell to maximize her weekly sales?
the inequalitites i got were:
s+l<(or equal to)15
8s+7l<(or equal to)585
s>(or equal to)0
l>(or equal to)0


is that correct?

Answer by stanbon(75887)

Mrs. Levy sells cookie bouquets. She can make no more than 15 bouquets in one week. Small bouquets sell for $25 and large bouquets sell for $55. If she prefers to make no more than 7 large arrangements a week, how many of each size should she sell to maximize her weekly sales?
--------------
0<=L<=7
0<=S<=15
S+L<=15
Graphing these inequalities on an S-L coordinate system you get
intersecting points at (0,0)(15,0)(8,7)(0,7)
Evaluating the S,L combinations of these points you get:
(0.0)=0
(15,0)=15
25=375 dollars
(8,7)= 8*25+7*55=585 dollars
(0,7)= 0*25+7*55=385 dollars
So the maximum comes from 8 small and 7 large bouquets.
cheers,
stan H.